Solve
(2x + 3y = 7
(x - 2y = 7 (Use substitution.)

Answer by nabla(475)

From the second equation,
(1) x=7+2y
plug (1) into the first equation:
2(7+2y)+3y=7
14+4y+3y=7
7y+14=7
(2) y=-1
Now, plug (2) into (1):
x=7+2(-1)
(3) x=5
Now, check to make sure these are correct:
2(5) + 3(-1) =10-3=7
5 - 2(-1) = 7